Given the Earnings before interest and taxes(EBIT) is Rs 30000 Interest payment is Rs 10000 Dividend on the preference shares Rs9000 Taxes being 50% of the profit before taxes(PBT). Number of outstanding equity shares 10000

What would be the earning per share(EPS) and degree of financial leverage (DFL), What would be the change in EPS, and DFL, if the EBIT increases to Rs50000 and Rs80000 Note- you have to calculate three EPS and three calculations for DFL


1
Expert's answer
2021-11-28T17:57:13-0500

Particulars Option 1 Option 2 Option 3

EBIT 30,000  50,000 80,000

(-) Interest 10,000 10,000  10,000

EBT 20,000 40,000 70,000

Tax @ 50% 10,000 20,000 35,000

EAT 10,000 20,000 35,000

Pref. Dividend 9,000 9,000 9,000

Earnings for equity 1,000 11,000  26,000

EPS = 0.1 1.1 2.6

EPS = (Earnings for equity/Number of equity shares)

DFL = 1.5 1.25 1.14

DFL = EBIT / EBT

Change in EPS =

When EBIT is 50,000 = (1.1 - 0.1) w 100 / 0.1 = 1000%,

When EBIT is 80,000 = (2.6 - 0.1) × 100 / 0.1 = 2,500%

Change in DFL =

When EBIT is 50,000 = (1.25 - 1.5) × 100 / 1.5 = -16.67%

When EBIT is 80,000 = (1.14 - 1.5) × 100 / 1.5 = -24%